# Is Your AI Truly Yours? Leveraging Blockchain for Copyrights, Provenance, and Lineage

### v1

@@ Abstract
As Artificial Intelligence (AI) integrates into diverse areas, particularly in content generation, ensuring rightful ownership and ethical use becomes paramount. AI service providers are expected to prioritize responsibly sourcing training data and obtaining licenses from data owners. However, existing studies primarily center on safeguarding static copyrights, which simply treats metadata/datasets as non-fungible items with transferable/trading capabilities, neglecting the dynamic nature of training procedures that can shape an ongoing trajectory.
In this paper, we present IBis, a blockchain-based framework tailored for AI model training workflows. IBis integrates on-chain registries for datasets, licenses and models, alongside off-chain signing services to facilitate collaboration among multiple participants. Our framework addresses concerns regarding data and model provenance and copyright compliance. IBis enables iterative model retraining and fine-tuning, and offers flexible license checks and renewals. Further, IBis provides APIs designed for seamless integration with existing contract management software, minimizing disruptions to established model training processes. We implement IBis using Daml on the Canton blockchain. Evaluation results showcase the feasibility and scalability of IBis across varying numbers of users, datasets, models, and licenses.
@@ I Introduction
The proliferation of Large Language Models (LLMs) based applications [1, 2] represents a significant milestone in the integration of Artificial Intelligence (AI) technologies into various facets of daily life, spanning from information retrieval [3, 4] to content generation [5, 6]. Concurrently, AI service providers have made strides in commercializing their services. Nevertheless, as LLMs and other AI models rely on extensive datasets aggregated from diverse sources for training [7, 8], apprehensions have emerged regarding the potential infringement of copyrights [9, 10, 11] during the data acquirement and model training process. To uphold responsible and ethical AI practices [12, 13], comply with regulations, and reduce legal liabilities, AI service providers must actively collaborate with data owners, including content creators and media industry stakeholders. Establishing licensing agreements [14, 15] and obtaining consent before utilizing data for AI model training is a key element of this collaboration [16]. Hence, there is a growing need for new frameworks addressing data provenance, lineage, and copyright compliance in the AI industry, tailored to its distinct needs and workflows.
@@
We have identified a series of functional challenges that must be addressed in the development of a such blockchain-based compliance framework: (i) The framework must be designed to seamlessly integrate with the existing workflow of AI model training. (ii) The framework should support continuous model retraining and fine-tuning with new datasets, allowing for the generation of updated models while maintaining data provenance and lineage. (iii) The framework should support mechanisms for license expiration and renewal, accommodating diverse business models employed by data owners. (iv) The ownership of datasets and models, along with all training actions, should be accompanied by evidence to clarify their licensing scope and ensure accountability for any subsequent actions. (v) The framework should facilitate communication between AI service providers and data owners, enabling efficient attainment and documentation of licensing agreements. (vi) The framework should ensure the effective management and commercial sensitivity of licenses, safeguarding them against unauthorized access by third parties.
@@
      Traceable registry (addressing c-iv): By deploying three on-chain, immutable registries for dataset metadata, licenses, and model metadata, the framework maintains authentic records of dataset and model relationships, ownership, and their copyright agreements. The bidirectional links between these records enables two-way traceability throughout data and model copyright management, provenance, and lineage processes.
@@ II-A Background
Copyrights. Copyright grants creators exclusive rights to their original expressions such as literary, artistic, and musical works. This legal framework safeguards creators’ rights, allowing them to control how their work is used, reproduced, and distributed.
Copyright protection is automatic upon creation, but it is implicit, requiring additional steps for proper protection. First, registering the work with the copyright office offers authoritative legal evidence of ownership and eligibility for statutory damages in case of infringement. Second, adding a copyright notice (©) with the creator’s name and the year of creation informs others of the copyright claim [30], akin to a signature on a picture. Additionally, NFTs [31] offer a novel method for embedding ownership of digital art via blockchain technology, with ownership automatically claimed upon minting.
@@ II-B Related Work
Protecting copyright/data in AI. Data and copyright protection in AI services is a long-standing topic. Existing methods can classified in several aspects [32]: Data-modifying approaches involve modifying or sanitizing user data to unlink them from specific individuals (e.g., k-anonymity [33], differential privacy [34], and watermarking [35]). This minimizes the risk of reidentification by removing or concealing Personally Identifiable Information (PII). Data-encrypting approaches encrypt user data to ensure integrity and confidentiality during data sharing, leveraging techniques such as homomorphic encryption [36] and secure Multi-Party Computation (MPC) [37, 38]. Data-minimizing approaches aim to boost efficiency by reducing the volume of personal data needed [39], often observed in general model training where PII data are not required during training and minimally during inference. Data-confining approaches involve AI methods that operate without sharing PII data beyond user boundaries [40], ensuring data integrity and confidentiality while enabling effective personalization through local access to personal data.
@@ III-A Design Overview
Architecture. We envision an ecosystem that empowers both AOs and COs to harness the network effect of a unified platform to train and use numerous AI models and datasets. For example, an CO could license the same dataset to multiple AOs and reap the benefits of a pay-as-you-go model for dataset usage or derived work within the same platform. Therefore, our proposed framework, IBis (cf. Fig.2), is designed to integrate a blockchain network hosted by a subset of AOs and COs. While established and commercially significant AOs and COs may operate blockchain nodes, others only require the capability to connect to one of them via an agent.
@@
Figure 2: Architecture design.
Key modules. IBis has the following six main modules (marked by green in Fig.2):
@@
  • License Validity Check (LVC) employs smart contracts to verify the validity of a license based on a set of environment variables, including the current date, AO’s operating location, and any other variables that could potentially contravene terms and conditions stipulated in the license. Our framework allows the creation of custom LVC smart contracts targeting different license types.
Stages. For the initial model training, the workflow of our framework can be segmented into the following three stages:
@@
Specifically, the workflow begins with dataset categorization, where datasets are organized into specific categories based on their content, source, and usage. This categorization facilitates efficient retrieval and management of datasets throughout the AI model development process. Following categorization, metadata registration takes place via DMR, recording crucial details such as data descriptions, authorship information, and usage rights within a structured format. This step ensures that comprehensive information about the datasets is readily accessible and referenced during their lifecycle. Finally, license checks are conducted via LVC smart contracts, utilizing automated processes to verify the authenticity and compliance of licenses associated with the datasets. Smart contracts ensure that AI model training and usage adhere to copyright agreements providing a streamlined and compliant workflow for managing datasets and their associated licenses.
@@
S3. Model metadata registering and copyright owner notification: In this stage, post-training models are recorded in on-chain MMR, creating a bidirectional linkage between models and training datasets. Notifications may be dispatched to COs as stipulated by the licensing agreement.
III-B Details of Each Stage
@@
III-B1 Dataset metadata registering and license check
@@
III-B2 License agreement drafting and bilateral signing
@@
III-B3 Model registration and CO notification
@@
III-C AI Model Retraining and Fine-tuning
@@
However, in retraining and fine-tuning scenarios, it is possible that the original model may need a license renewal at the time of retraining or fine-tuning, as one or more licenses of its training datasets might have become invalid (e.g., expired). Therefore, before retraining or fine-tuning the model, an additional stage is necessary to verify data eligibility by determining if the model requires a license renewal. Fig.3(b) depicts the flowchart of actions during model retraining or fine-tuning. Sec.IV-C1 presents the detailed mechanism used to conduct the license renewal check.
@@
IV Detailed Construction
@@ IV-A Data Models
The main attributes that IBis framework supports can be broadly grouped as follows (see Table I):
@@
TABLE I: Dataset, license, and model attributes.
@@
Dataset $\mathsf{datasetId}$
Unique identifier of the dataset.
$\mathsf{sourceUrl}$
URL from which the dataset was scraped.
$\mathsf{copyrightOwnerId}$
Unique identifier of the copyright owner.
$\mathsf{licenseId}$
Unique identifier of the copyright license.
$\mathsf{modelList}$
List of unique identifiers of the models trained on this dataset.
$\mathsf{modelOwnerId}$
Unique identifier of AO who scrapes and adds this dataset.
License $\mathsf{licenseId}$
Unique identifier of the dataset.
$\mathsf{scope}$
URL to dataset. Datasets pointed by this URL fall within the scope of the license.
$\mathsf{copyrightOwnerId}$
Unique identifier of the copyright owner.
$\mathsf{copyrightOwnerSignature}$
Digital signature of the copyright owner.
$\mathsf{modelOwnerId}$
Unique identifier of the model owner.
$\mathsf{modelOwnerSignature}$
Digital signature of the model owner.
$\mathsf{validFrom}$
Timestamp indicating when the license takes effect.
$\mathsf{typeId}$
Unique identifier of the license type. Used to determine the smart contract to check the license validity.
$\mathsf{datasetList}$
List of identifiers of the datasets covered by this license.
Model $\mathsf{modelId}$
Unique identifier of the AI model.
$\mathsf{modelOwnerId}$
Unique identifier of the model owner.
$\mathsf{datasetList}$
List of unique identifiers of training datasets.
$\mathsf{sourceModelId}$
Unique identifier of the source model that has been retrained.
$\mathsf{childModelList}$
List of identifiers of the models trained based on this model.
We highlight two aspects. First, the data model is extensible, enabling AOs and COs to incorporate additional custom attributes as needed. For example, a storage URL can be included in dataset metadata to indicate where AO stores the dataset. License can include custom attributes such as expiration date, exclusivity, and other terms and conditions. Second, a web path is employed to delineate the scope of what is being licensed, considering that the majority of AI models are trained using online data.
@@
IV-B Functional Operations
This section delineates the operations that can be performed by AOs and COs, along with their time complexity analysis. Table II lists the time complexity of operations and the entities authorized to perform them. We assume that the on-chain license registry, DMR, and MMR are implemented as hash maps on a smart contract, resulting in a time complexity of $O(1)$ for searching them.
Figure 6: Model, dataset, and license relationships.
@@ IV-B Functional Operations
  • Search for a relevant license: This operation extracts the dataset’s $\mathsf{copyrightOwnerId}$ and performs a search on the license registry using $\mathsf{copyrightOwnerId}$, which involves a complexity of $O(1)$. This search may yield a list of licenses with the same $\mathsf{copyrightOwnerId}$ (albeit with different scopes). Finally, a scan is conducted on the list of licenses to filter out the licenses with irrelevant scopes. Our framework operates on the premise that license scopes do not intersect, ensuring each dataset corresponds to at most one license. As CO is unlikely to have many licenses with the same AO for practical reasons, one can assume the size of this list to be small. Consequently, we can still assume the overall time complexity to be $O(1)$.
TABLE II: Operations, complexities, and authorizations.
Operation Complexity Authorisations
$\mathsf{getDatasetLicense}$ $O(1)$ AOs
$\mathsf{getModelLicenses}$ $O(|D|+|M|)$ AOs
$\mathsf{checkLicenseValidity}$ $O(|E|)$ AOs/COs
$\mathsf{getLicensedDatasets}$ $O(|D||E|)$ AOs
$\mathsf{getDatasetsByLicense}$ $O(1)$ AOs
$\mathsf{getModelsByLicense}$ $O(|D||M|)$ AOs
$\mathsf{getModelDatasets}$ $O(|M|)$ AOs
Obtain model license. Given a model identifier $\mathsf{modelId}$, the $\mathsf{getModelLicenses}$ operation retrieves the licenses of its training datasets. This operation requires executing $\mathsf{getDatasetLicense}$ for each of xthe training datasets of the provided model and its upstream source models. To identify the contributing training datasets, the operation functions as a graph traversal algorithm on a graph with the given model as the root node, the upstream models as intermediate nodes, and the datasets as leaf nodes (as depicted in Fig.6). The time complexity of a basic graph traversal algorithm is $O(|V|+|E|)$, where V is the set of vertices and E is the set of edges. Therefore, given a graph with a set of M models and D datasets, the time complexity of the graph traversal is $O(|D|+|M|)$. As $\mathsf{getDatasetLicense}$’s complexity is $O(1)$, the graph traversal dominates the overall time complexity.

### v2

@@ Abstract
As Artificial Intelligence (AI) integrates into diverse areas, particularly in content generation, ensuring rightful ownership and ethical use becomes paramount, AI service providers are expected to prioritize responsibly sourcing training data and obtaining licenses from data owners. However, existing studies primarily center on safeguarding static copyrights, which simply treat metadata/datasets as non-fungible items with transferable/trading capabilities, neglecting the dynamic nature of training procedures that can shape an ongoing trajectory. In this paper, we present IBis, a blockchain-based framework tailored for AI model training workflows. Our design can dynamically manage copyright compliance and data provenance in decentralized AI model training processes, ensuring that intellectual property rights are respected throughout iterative model enhancements and licensing updates. Technically, IBis integrates on-chain registries for datasets, licenses and models, alongside off-chain signing services to facilitate collaboration among multiple participants. Further, IBis provides APIs designed for seamless integration with existing contract management software, minimizing disruptions to established model training processes. We implement IBis using Daml on the Canton blockchain. Evaluation results showcase the feasibility and scalability of IBis across varying numbers of users, datasets, models, and licenses.
@@ I Introduction
Artificial Intelligence (AI) technologies [1, 2] have increasingly permeated various aspects of daily life, spanning from information retrieval [3, 4] to content generation [5, 6]. Concurrently, AI service providers have made strides in commercializing their services. Nevertheless, as AI models rely on extensive datasets aggregated from diverse sources for training [7, 8], apprehensions have emerged regarding the potential infringement of copyrights [9, 10, 11] during the data acquirement and model training process. To uphold responsible and ethical AI practices [12, 13], comply with regulations, and reduce legal liabilities, AI service providers must actively collaborate with data owners, including content creators and media industry stakeholders. Establishing licensing agreements [14, 15] and obtaining consent before utilizing data for AI model training is a key element of this collaboration [16]. Hence, there is a growing need for new frameworks addressing data provenance, lineage, and copyright compliance in the AI industry, tailored to its distinct needs and workflows.
@@
We have identified a series of functional challenges for such a blockchain-based compliance framework [27, 28, 29]: (i) The framework needs to be designed to seamlessly integrate with the existing workflow of AI model training [30, 31, 32]. (ii) The framework should support continuous model retraining and fine-tuning with new datasets [33], allowing for the generation of updated models while maintaining data provenance and lineage. (iii) The framework should support mechanisms for license expiration and renewal, accommodating diverse business models employed by data owners. (iv) The ownership of datasets and models [34, 35], along with all training actions, should be accompanied by evidence to clarify their licensing scope and ensure accountability for any subsequent actions. (v) The framework should facilitate communication between AI service providers and data owners [19, 36], enabling efficient attainment and documentation of licensing agreements. (vi) The framework should ensure the effective management and commercial sensitivity of licenses, safeguarding them against unauthorized access by third parties [37].
@@
      Traceable registry (addressing c-iv): By deploying the on-chain, immutable registries for dataset metadata, licenses, and model metadata, the framework maintains authentic records of dataset and model relationships, ownership, and their copyright agreements. The bidirectional links between these records enables two-way traceability throughout data and model copyright management, provenance, and lineage processes.
@@ II-A Background
Copyrights. Copyright grants creators exclusive rights to their original expressions, such as literary, artistic, and musical works. This legal framework safeguards creators’ rights, allowing them to control how their work is used, reproduced, and distributed.
Copyright protection is automatic upon creation, but it is implicit, requiring additional steps for proper protection. First, registering the work with the copyright office offers authoritative legal evidence of ownership and eligibility for statutory damages in case of infringement. Second, adding a copyright notice (©) with the creator’s name and the year of creation informs others of the copyright claim [41], akin to a signature on a picture. Additionally, NFTs [42] offer a novel method for embedding ownership of digital art via blockchain, with ownership automatically claimed upon minting.
@@
TABLE I: Comparison of IBis with related solutions
Solutions c-i: Traceability c-ii: Immutability c-iii: Multi-party Licensing c-iv: Reuse Control c-v: Lineage c-vi: Scalable Querying
Generic NFTs [35] $\times$ $\times$ $\times$ $\times$
Provenance DRM [43, 44, 45] $\times$ $\times$ $\times$
LicenseChain [46, 47, 48, 49] $\sim$ $\times$ $\times$
IBis (this work)
@@ II-B Related Work
Protecting copyright/data in AI. Data and copyright protection in AI services is a long-standing topic. Existing methods can be classified in several aspects [50]: Data-modifying approaches involve modifying or sanitizing user data to unlink them from specific individuals (e.g., k-anonymity [51], differential privacy [52], and watermarking [53]). This minimizes the risk of reidentification by removing or concealing Personally Identifiable Information (PII). Data-encrypting approaches encrypt user data to ensure integrity and confidentiality during data sharing, leveraging techniques such as homomorphic encryption [54] and secure Multi-Party Computation (MPC) [55]. Data-minimizing approaches aim to boost efficiency by reducing the volume of personal data needed [56], often observed in general model training where PII data are not required during training and minimally during inference. Data-confining approaches involve AI methods that operate without sharing PII data beyond user boundaries [57], ensuring data integrity and confidentiality while enabling effective personalization through local access to personal data.
@@
Comparison. Table I summarizes how IBis compares with representative blockchain-based copyright, provenance, and licensing systems. We align our comparison with the six core challenges (c-i to c-vi) outlined in the Introduction. Existing platforms often address immutability or traceability, but fall short in fine-grained license control, model retraining lineage, and scalable graph-based querying.
Our design diverges from these existing models by opting not to rely on computationally heavy cryptographic primitives. Instead, we utilize blockchain-based multi-party signing and customizable access control mechanisms, which streamline the establishment and renewal of licensing agreements. This ensures secure and transparent lineage tracking directly within AI model training workflows, offering a novel approach that enhances both compliance and operational efficiency without overhead associated with traditional cryptographic methods.
@@ III-A Design Overview
Architecture. We envision an ecosystem that empowers both AOs and COs to harness the network effect of a unified platform to train and use numerous AI models and datasets. For example, CO could license the same dataset to multiple AOs and reap the benefits of a pay-as-you-go model for dataset usage or derived work within the same platform. Therefore, our proposed framework, IBis (cf. Fig.2), is designed to integrate a blockchain network hosted by a subset of AOs and COs. While established and commercially significant AOs and COs may operate blockchain nodes, others only require the capability to connect to one of them via an agent.
@@
Figure 2: Architectural design.
III-B Key Modules
IBis has the following six main modules (green in Fig.2):
@@ III-B Key Modules
  • License Validity Check (LVC) employs smart contracts to verify the validity of a license based on a set of environment variables, including the current date, AO’s operating location, and any other variables that could potentially contravene terms and conditions stipulated in the license. Our framework allows the creation of custom LVC smart contracts targeting different licence types.
Notably, both MSS and CLM are implemented as off-chain services. They are essential for facilitating interactions that precede and follow blockchain transactions, but they do not operate directly on the blockchain.
III-C Stages
For the initial model training, the workflow of our framework can be segmented into the following three stages:
@@ III-C Stages
Specifically, the workflow begins with dataset categorization, where datasets are organized into specific categories based on their content, source, and usage. This categorization facilitates efficient retrieval and management of datasets throughout the AI model development process. Following categorization, metadata registration takes place via DMR, recording crucial details such as data descriptions, authorship information, and usage rights within a structured format. This step ensures that comprehensive information about the datasets is readily accessible and referenced during their lifecycle. Finally, license checks are conducted via LVC smart contracts, utilizing automated processes to verify the authenticity and compliance of licenses associated with the datasets. Smart contracts ensure that AI model training and usage adhere to copyright agreements, providing a streamlined and compliant workflow for managing datasets and their associated licenses.
@@
S3. Model registering and copyright owner notification: In this stage, post-training models are recorded in on-chain MMR, creating a bidirectional linkage between models and training datasets. Notifications may be dispatched to COs as stipulated by the licensing agreement.
III-D Details of Each Stage
@@
III-D1 Dataset metadata registering and license check
@@
III-D2 License agreement drafting and bilateral signing
@@
III-D3 Model registration and CO notification
@@
III-E AI Model Retraining and Fine-tuning
@@
However, in retraining and fine-tuning scenarios, it is possible that the original model may need a license renewal at the time of retraining or fine-tuning, as one or more licenses of its training datasets might have become invalid (e.g., expired). Therefore, before retraining or fine-tuning the model, an additional stage is necessary to verify data eligibility by determining if the model requires a license renewal. Fig.3(b) depicts the flowchart of actions during model retraining or fine-tuning. Sec.IV-E presents the detailed mechanism used to conduct the license renewal check.
@@
IV IBIS with Detailed Workflow
@@ IV-A Data Models
The main attributes that IBis framework supports can be broadly grouped as follows (see Table II):
@@
TABLE II: Dataset, license, and model attributes.
@@
Dataset $\mathsf{datasetId}$ Unique identifier of the dataset.
$\mathsf{sourceUrl}$ URL from which the dataset was scraped.
$\mathsf{copyrightOwnerId}$ Unique identifier of the copyright owner.
$\mathsf{licenseId}$ Unique identifier of the copyright license.
$\mathsf{modelList}$ List of unique identifiers of the models trained on this dataset.
$\mathsf{modelOwnerId}$ Unique identifier of AO who scrapes and adds this dataset.
License $\mathsf{licenseId}$ Unique identifier of the dataset.
$\mathsf{scope}$ URL to dataset. Datasets pointed by this URL fall within the scope of the license.
$\mathsf{copyrightOwnerId}$ Unique identifier of the copyright owner.
$\mathsf{copyrightOwnerSignature}$ Digital signature of the copyright owner.
$\mathsf{modelOwnerId}$ Unique identifier of the model owner.
$\mathsf{modelOwnerSignature}$ Digital signature of the model owner.
$\mathsf{validFrom}$ Timestamp indicating when the license takes effect.
$\mathsf{typeId}$ Unique identifier of the license type. Used to determine the smart contract to check the license validity.
$\mathsf{datasetList}$ List of identifiers of the datasets covered by this license.
Model $\mathsf{modelId}$ Unique identifier of the AI model.
$\mathsf{modelOwnerId}$ Unique identifier of the model owner.
$\mathsf{datasetList}$ List of unique identifiers of training datasets.
$\mathsf{sourceModelId}$ Unique identifier of the source model that has been retrained.
$\mathsf{childModelList}$ List of identifiers of the models trained based on this model.
We highlight two aspects. First, the data model is extensible, enabling AOs and COs to incorporate additional custom attributes as needed. For example, a storage URL can be included in dataset
Figure 6: Model, dataset, and license relationships.
metadata to indicate where AO stores the dataset. A license can include custom attributes such as expiration date, exclusivity, and other terms and conditions. Second, a web path is employed to delineate the scope of what is being licensed, considering that the majority of AI models are trained using online data.
@@
Integrity binding with off-chain data. While the actual datasets and models are stored off-chain, IBis ensures integrity through the use of cryptographic hashes. Each dataset and model is associated with a content-addressable identifier, typically computed using a collision-resistant hash function (e.g., SHA-256). These hashes are recorded on-chain as part of the metadata entries in the DMR and MMR smart contracts. As a result, any tampering with off-chain content can be detected by recomputing and comparing the hash values.
Optionally, the system may integrate with decentralized storage platforms (e.g., IPFS), where the content hash also serves as the retrieval address. This dual use of hashes enhances verifiability and offers a robust mechanism for ensuring data provenance and integrity even in off-chain storage.
IV-B Obtaining Licenses
@@
  • Search for a relevant license: This operation extracts the dataset’s $\mathsf{copyrightOwnerId}$ and performs a search on the license registry using $\mathsf{copyrightOwnerId}$, which involves a complexity of $O(1)$. This search may yield a list of licenses with the same $\mathsf{copyrightOwnerId}$ (albeit with different scopes). Finally, a scan is conducted on the list of licenses to filter out the licenses with irrelevant scopes. Our framework operates on the premise that license scopes do not intersect, ensuring each dataset corresponds to at most one license. As CO is unlikely to have many licenses with the same AO for practical reasons, one can assume the size of this list to be small.
Obtain model licenses. Given a model identifier $\mathsf{modelId}$, the $\mathsf{getModelLicenses}$ operation retrieves the licenses of its training datasets. This operation begins by identifying the training datasets associated with the provided model and its upstream source models. Each dataset is linked to a single license, and the operation retrieves these licenses by traversing the graph where models, datasets, and licenses are nodes. The graph includes edges representing model-to-dataset, dataset-to-license, and model-to-model relationships (as depicted in Fig.6). For a specific model, the operation traverses $T$ datasets linked to it, retrieves their associated licenses, and deduplicates them if necessary. The time complexity for this operation is $O(T)$ per model, as retrieving the licenses for $T$ datasets involves processing $T$ edges. When the operation considers all $M$ models, the total complexity becomes $O(|M||T|)$, where $T$ is the number of datasets linked to each model. This complexity accounts for accessing training datasets, retrieving licenses, and aggregating results. Since $T$ is typically much smaller than $D$ (the total number of datasets in the graph), the traversal focuses only on the relevant portion of the graph. Therefore, the overall complexity scales linearly with the number of models and their associated datasets.
IV-C Validating Licenses